If Mary's iphone charger is 6 1/2 feet long and her bed is 9 feet from the outlet in the wall, how much longer would her charger need to be for her to be able to charge her phone in bed?



1
Expert's answer
2020-10-05T16:40:08-0400

Solution. To charge an iPhone in bed, the charger must be 9 feet long. Therefore, the charger must be longer by


"9-6\frac{1}{2}=2\frac{1}{2} feet"

Answer.

"2\frac{1}{2} feet"